Grade 11Mechanics

A body starts from rest with constant acceleration its velocity after 2 n seconds is v find displacement of body in last n seconds

Bhawika
9 Years agoGrade 11

1 Answer

Umakant biswal
9 Years ago
@ BHAWIKA 
i am giving the approach to solve the problem , follow that 
here , u=0, 
velocity after 2n sec, will be given as v=u+at=0+a*2n. 
so, acceleration, a=v/2n
so, velocity after n sec..- 
v1=0+at=(v/2n)*n=v/2
now, apply eq. of motion to find d in last n sec 
v^2-v1^2=2ad
put values of v,v1 and a
d=[v^2-(v/2)^2]/2*(v/2n)=(3/4)v/(v/n)=3/4*n
i think u can solve it right now . 
if not then let me know in this post , i will solve the question fully . 
ALL THE BEST ..